The current drawn from the cell when the wire is connected across A as shown is 0.3A but reduced to 0.25A, when the wire is removed leaving



A. Alone connected to the the cell determine



B. The E.M.F of the cell



C. The internal resistance of the cell

1
Expert's answer
2022-03-23T14:00:16-0400

(1)


(B)

E=V-ir

"E=V-0.3r"

E=V-0.25r

(C)V=E-ir

r=Internal resistance